OpenStudy (anonymous):

Sergio purchased 4 items for $1.50 each. sales tax was 8.25%. what was the total price?

OpenStudy (jdoe0001):

\(\bf 4\times 1.50\implies \square \\ \quad \\ {\color{brown}{ 8.25}}\%\ of\ \square \implies \cfrac{{\color{brown}{ 8.25}}}{100}\cdot \square \implies {\color{brown}{ 8.25}}\div 100\cdot \square \\ \quad \\ \textit{total price is }\implies \square +({\color{brown}{ 8.25}}\div 100\cdot \square)\)